# Zhiwei Zuo Zhuo Tang Kenli Li Anwitaman Datta

### v2

@@ 3.2 Our Approach
The parameters in the machine learning model store the high-dimensional features of the data, so in order to eliminate the influence of revoked data on the model, inspired by adversary learning szegedy2013intriguing , we consider adding slight perturbation to trained model and then perform several training epochs. Current machine unlearning methods based on parameters perturbation basically change all the parameters of the model or use tricks such as regularization or dropout to improve the efficiency of unlearning. We desire to perturb the parameters of the model at a finer granularity.
@@
Not all parameters affect the model equally. The impact of the former layers on the model is mainly to extract the features of the data, while the later layers use these features to perform tasks. To this extent, it can be considered that perturbing the last few layers can achieve the purpose of inexact unlearning. However, only modifying the last few layers also means a lower degree of machine unlearning, as shown in Tab.2, where the accuracy $Acc_{UL}$ on $D_{UL}$ for EU-5, EU-10, CF-5, CF-10 is significantly higher than other unlearning strategies. In order to make the perturbation ratio decrease and make the perturbation more precise, we consider identifying several parameters in the model that have the greatest impact on performance to perturb, or to address the issue of reduced unlearning effect caused by perturbing the parameters of the last few layers. Thus, we propose two strategies for perturbing model parameters, namely Random-k and Top-K.
3.2.1 Parameter Selection and Perturbation
@@
For the Top-K strategy, as previously mentioned, it is necessary to evaluate the sensitivity of each model parameter. Certainly, perturbing model parameters one by one can determine their sensitivity, but it’s time-consuming. To quickly obtain the sensitivity of parameters, we approximate it using the gradient. Specifically, after the model has converged, yielding the function $F(x,w)$, we introduce a small perturbation to the model parameters denoted as $\epsilon=\{\epsilon_{i,j}\}$. The change in the model’s output can be approximately measured using the following formula:
@@
Where $\bigtriangleup(x,w)$ represents the change in the model’s output due to the introduced perturbation $\epsilon$, and $\frac{\partial F(x,w)}{\partial\theta_{i,j}}=g_{i,j}(x)$ is the gradient of learned model to parameter $w$. This evaluation of sensitivity allows us to assess the impact of each parameter on the model’s predictions.
Focusing on aforementioned Equation 3, assuming the perturbation $\epsilon$ added to the model is the same constant, the calculation of sensitivity can be simplified as follows:
@@
Fig.2 illustrates the process of our approach. After completing the parameter screening, we introduce perturbations by drawing upon the insights from szegedy2013intriguing . Specifically, we apply a minute amount of noise to the selected parameters as follows:
$w_{i}=\epsilon*w_{rand},\quad w_{i}\in\{Top-K(w)\lor{Random-k(w)}\}$ (5)
@@ 3.2.2 Unlearning Process
For the purpose of eliminating the influence of the data to be unlearned on the model, retraining from scratch, despite consuming substantial computational resources and time, can achieve the optimal unlearning effect. As far as the strategy for machine unlearning is concerned, we desire that the unlearning model $M_{UL}$ approximates the retraining model $M_{RE}$, thus ensuring that the unlearning is moving in the right direction. We aim for both the retraining model and the unlearning model to display similar performance when subjected to $D_{UL}$. Therefore, in the third step of unlearning, we input the data intended for unlearning into both models to obtain the corresponding output distributions.
Jensen-Shannon (JS) divergence is used to guide the model training of the unlearning process chundawat2023can . As a symmetric and always finite measure, JS divergence provides a reliable measure of the similarity or distance between two probability distributions. Suppose $x\in D_{UL}$, $P(w,x)$ and $P(\theta,x)$ are posterior distributions of retraining model $M_{RE}$ and unlearning model $M_{UL}$, respectively. $JS$ divergence is calculated as following:
@@
$JS(P(\theta,x)||P(w,x))=\frac{1}{2}\sum P(\theta,x)\log P(\theta,x)+$ (7)
$\begin{split}\frac{1}{2}\sum P(w,x)\log P(w,x)-\log(P(\theta,x)+P(w,x))-\log 2\end{split}$
In the context of machine unlearning, the JS divergence between distributions $P(w,x)$ and $P(\theta,x)$ can quantify the influence of $D_{UL}$ on $M_{UL}$. If the JS divergence is small, the influence of $D_{UL}$ on $M_{UL}$ is deemed to be minor. Therefore, during training, our goal should be to minimize the JS divergence between these distributions to ensure the effectiveness of unlearning. Moreover, given that the value of the JS divergence always falls between 0 and 1, we can intuitively understand the degree of similarity between the two distributions.
@@
8:     $w_{j}\leftarrow\epsilon*w_{rand},w_{j}\in\{Top-K(w)\}$
@@
11:     $mask\leftarrow(mask<k)*\epsilon$
12:     $w\leftarrow(1-mask)*w+mask*w_{rand}$
@@
3.3 Unlearning Effectiveness
By employing a fine-grained model parameter perturbation strategy for the machine unlearning process, the unlearning model and the retraining model essentially exhibit the characteristic of model indistinguishability. For datasets with i.i.d attributes, quantifying the unlearning effect by examining the accuracy on the remaining data $D_{RE}$ and the unlearning data $D_{UL}$ of the unlearning model can be challenging. As Fig.1 and Fig.6 show (accompanied with detailed discussions of results later), the accuracy of both $D_{RE}$ and $D_{UL}$ may be maintained at a fairly satisfactory level, the accuracy difference between $D_{RE}$ and $D_{UL}$ is quite small.
Figure 3: SPD-GAN architecture
To address the difficulty in measuring the effect of unlearning and further quantify the degree of forgetting caused by each unlearning strategies, we consider breaking the i.i.d attribute of $D_{UL}$.
@@ 3.3 Unlearning Effectiveness
where $\eta$ is a coefficient to balance the $M$’s classification accuracy and $M_{UL}$’s unlearning effect. As Equation.1 explained, our goal is to explicitly demonstrate the degree of unlearning in terms of performance. Once completing the training of SPD-GAN, the performance difference between $M(\cdot)$ and $M_{UL}(\cdot)$ on perturbed data $D_{p}$ is indicative of unlearning degree. That is:
@@
3.4 Time Complexity Analysis
@@
We use Equation 4 to approximate the sensitivity of each parameter. Since $|w|=N$, the time complexity for both the forward pass and back-propagation is linear in the number of parameters, the overall time complexity for evaluating sensitivity remains $\mathcal{O}(N)$. Once the model structure is determined, the parameters with Top-K sensitivity of a specific model should be calculated only once.
4 Experiments
4.1 Baseline Methods
The impact of the first few layers on the model is mainly to extract the features of the data, and the latter few layers use these features to perform tasks. Following the setting of DBLP:journals/corr/abs-2201-06640 , we set $K=\{5,10\}$, use EU-K and CF-K as two of our baseline methods.
EU-K: Performing exact unlearning involves targeting the final $K$ layers of the model. In this process, the model parameters are held constant except the last $K$ layers. These last $K$ layers are then reinitialized and trained from the beginning.
CF-K: Catastrophic forgetting the last $K$ layers of the model. Differing from EU-K, there’s no need to reinitialize the parameters of the last $K$ layers. The parameters of all layers except the last $K$ are held constant.
@@ 4.2 Evaluation Metrics
The forgetting rate is a measure of performance decay on the unlearning data, post the model’s unlearning phase. In essence, the forgetting rate quantifies the extent of data ‘forgotten’ by the model:
@@
Here, $Acc_{before}$ represents the model’s accuracy on the unlearning data prior to the unlearning process, and $Acc_{after}$ indicates the model’s accuracy on the same set, post-unlearning.
@@
To assess the degree of model indistinguishability, we employ the $similarity=1-JS~{}divergence$ as a metric to quantify the similarity between the unlearning model and the retraining model.
4.3 Experimental Results
4.3.1 Training settings
We conduct experiments on CIFAR-10 krizhevsky2009learning datasets using ResNet18 he2016deep , VGG simonyan2014very , GoogLeNet szegedy2015going , and DenseNet huang2017densely models. The training dataset is randomly divided into $D_{UL}$ and $D_{RE}$ with $D_{UL}$ comprising 5%, 10%, 15% and 20%. For Top-K, we set $K=45$ for ResNet18, and for the Random-k, $k=5\%$. The coefficient $\lambda$ in Equation.8 is set to 0.1. We use Adam as optimizer, for EU-K and CF-K, we set learning rate 1e-4. The generator architecture in SPD-GAN consists of (two Conv2d)-BatchNorm-Conv2d-BatchNorm layers for the encoder, followed by ConvTranspose2d-BatchNorm-ConvTranspose2d-Conv2d layers for the decoder. The LeakyReLU and ReLU activation functions are used for the encoder and decoder, respectively. Activation function Tanh is used to limit the output within [-1, 1]. The coefficient $\eta$ in Equation.11 is 0.03. All the experiments are conducted on a Nvidia A100 GPU.
4.3.2 Determination of K value for Top-K
@@
To investigate the optimal value of $K$ for Top-K unlearning strategy, we conduct experiments with different values of $K$ within the range of 0 to 50 when the unlearning data ratio varies from 5% to 20%. Following the above settings, we apply Top-K to four models over 100 epochs. Since machine unlearning aims to eliminate the influence of unlearning data while trying to not hurt the performance of remaining data, we calculate the accuracy difference between $Acc_{RE}$ and $Acc_{UL}$. As illustrated in Fig.5, the $K$ value corresponding to the maximum accuracy difference is identified as the optimal. Considering the optimal K values under four unlearning data ratios, we set $K=45$ for ResNet18 and DenseNet, $K=30$ for VGG, and $K=50$ for GoogLeNet.
As shown in Fig.5, although we have chosen the $K$ corresponding to the maximum accuracy difference, in actual experiments, the selection of $K$ values fluctuating within the range of 0 to 50 does not bring significant variability to the Top-K unlearning. For example, in Fig.5(a), for ResNet18, setting K=25 versus 45 does not make a substantial difference. Therefore, one can avoid additional effort and computation required to determine any optimal $K$.
4.3.3 Effectiveness of Top-K and Random-k
Fig.6 illustrates the unlearning accuracy $Acc_{UL}$ and remaining accuracy $Acc_{RE}$ of four different models trained on the CIFAR-10 using Top-K and Random-k over 50 epochs with unlearning ratio 20%. We employ $Acc_{RE}$ instead of $Acc_{test}$ for two reasons. Firstly, we aim to account for the accuracy on $D_{RE}$. Secondly, the testing data also incorporates information closely related to $D_{UL}$, leading to a decrease in $Acc_{test}$. Fig.1 further shows the accuracy after applying the Random-k on the CIFAR-10 for 50 epochs. The accuracy of $D_{RE}$ is significantly higher than that of $D_{UL}$. It can be observed that the application of the Top-K achieves a 95% $Acc_{RE}$ across the four models, indicating a strong memory retention capability. Meanwhile, the accuracy difference between unlearning model on the $D_{RE}$ and the $D_{UL}$ shows that both the Top-K and Random-k achieve the unlearning effect more or less.
@@ 4.3.3 Effectiveness of Top-K and Random-k
Table 1: The number of parameters perturbed in different models under various unlearning strategies.
Ratio Strategy $Acc_{UL}$ $Acc_{RE}$ FR MRR Similarity Time
5% Top-K 80.01 97.61 19.97 97.63 88.37 76.97
Random-k 76.56 90.07 23.44 90.07 78.77 135.43
EU-5 85.68 98.44 14.31 98.44 84.88 534.76
EU-10 85.88 99.02 14.12 99.02 69.42 662.34
CF-5 86.52 99.16 13.48 99.16 78.65 586.6
CF-10 86.64 99.54 13.36 99.54 99.98 607.24
Retrain 84.84 100 15.16 100 100 1190.15
10% Top-K 80.14 97.19 19.86 97.19 85.36 192.86
Random-k 81.48 96.52 18.52 96.52 91.03 249.11
EU-5 86.7 98.88 13.29 98.88 82.25 541.38
EU-10 84.88 98.12 15.12 98.12 90.4 417.87
CF-5 85.52 97.89 14.48 97.89 70.27 515.13
CF-10 86.26 98.91 13.73 98.91 91.37 664.39
Retrain 84.8 100 15.2 100 100 1040.08
15% Top-K 75.48 96.8 24.52 96.8 89.88 203.23
Random-k 79.85 98.09 20.15 98.09 83.59 335.87
EU-5 84.61 97.46 15.38 97.46 85.11 494.60
EU-10 86.52 98.75 13.48 98.75 91.38 627.94
CF-5 87.64 99.46 12.35 99.46 85.27 558.31
CF-10 86.05 98.79 13.94 98.79 91.75 685.18
Retrain 84.85 100 15.15 100 100 950.51
20% Top-K 75.46 97.07 24.54 97.07 86.37 425.01
Random-k 77.83 98.99 22.17 98.99 94.51 638.63
EU-5 84.48 97.91 15.51 97.9 71.77 636.77
EU-10 83.19 96.81 16.81 96.81 83.92 741.51
CF-5 85.86 98.14 14.14 98.14 81.76 778.28
CF-10 85.69 99.17 14.31 99.17 91.9 979.65
Retrain 85.24 100 14.76 100 100 1400.9
@@
Table 2: The experimental results of the ResNet18 under different unlearning ratios of the CIFAR-10. Lower $Acc_{UL}(\%)$ and Time (s) and higher other metrics (%) indicate better performance.
@@
Tab.2 provides a more comprehensive overview of the effects of different unlearning strategies at various unlearning ratios. Across different unlearning ratios, Top-K consistently exhibits the lowest $Acc_{UL}$ values (80.14%, 75.48%, and 75.46%), along with corresponding forgetting rates (19.86%, 24.52%, and 24.54%). This suggests that, relative to other strategies, Top-K achieves the most effective reduction of the impact of $D_{UL}$ on the model when the unlearning ratio belongs to 10%, 15%, 20%. However, for the 5% unlearning ratio under Random-k, its corresponding $Acc_{RE}$ and MRR of 90.07% noticeably lag behind the results of other strategies. We believe this disparity is due to the instability of the strategy that employs random parameter perturbation, leading to inconsistent unlearning effects. In the meanwhile, across both $Acc_{RE}$ and MRR, the performance of $M_{UL}$ from nearly all unlearning strategies is comparable to that of retraining models, achieving close to 100% accuracy. As for similarity, there isn’t a significant distinction among different strategies, but this does indicate that model indistinguishability is maintained across diverse strategies.
(a) ResNet18
(b) VGG
(c) DenseNet
(d) GoogLeNet
@@
As discussed earlier, we believe that employing machine unlearning strategies offers a faster way to eliminate the influence of unlearning data on models compared to the retraining approach. The metric ’Time’ in Tab.2 indicates the time required for various strategies to complete unlearning. All strategies achieve an accelerations, but Top-K and Random-k take less time. As illustrated in Tab.1, this acceleration is due to the fact that, for experiments on ResNet18, Top-K and Random-k perturb less parameters than others. To further illustrate the acceleration effect of unlearning strategies on different models compared to the retraining method, we also conduct machine unlearning tasks on VGG, DenseNet, and GoogLeNet. Fig.7 shows the acceleration ratio of each unlearning strategy compared to retraining under different models and unlearning ratios. The acceleration effect of Top-K and Random-k is better than other strategies. Taking Fig.7(a) as an example, when the unlearning ratio is 5%, Top-K can achieve an acceleration of more than 15x, while Random-k achieves an acceleration close to 9x. The acceleration from EU-K and CF-K methods is not significant. As the unlearning ratio increases, the acceleration of Top-K and Random-k slows down. This is because, as more data is to be forgotten, the degree of modification to the model deepens. In conclusion, we can ascertain that the Top-K achieves optimal unlearning performance with minimal parameter perturbed, maintains considerable memory retention capability, and achieves the fastest acceleration effect.

### v3

@@ III-B Our Unlearning Approachs
The parameters in the machine learning model store the high-dimensional features of the data, so in order to eliminate the influence of unlearning data on the model, inspired by adversarial learning [36], we consider adding slight perturbation to trained model and then perform several fine-tuning epochs. Current machine unlearning methods based on parameters perturbation basically change all the parameters of the model or use tricks such as regularization or Dropout to improve the efficiency of unlearning. We desire to perturb the parameters of the model at a finer granularity.
Figure 3: SPD-GAN architecture
@@
Not all parameters affect the model equally. For instance, for images, the impact of the former layers on the model is mainly to extract the features (like color, shape, etc) of the data, while the later layers obtain high-dimensional information, thereby completing tasks such as object recognition. To this extent, it can be considered that perturbing the last few layers and freezing the former layers can achieve the purpose of inexact machine unlearning. However, only modifying the last few layers leads to a lower degree of machine unlearning, as shown in Tab.II, where the accuracy $Acc_{UL}$ on $D_{UL}$ for EU-5, EU-10, CF-5, CF-10 is significantly higher than other unlearning strategies.
In order to make the perturbation ratio decrease and make the perturbation more precise, we consider identifying several parameters in the model that have the greatest impact on performance (which we call ’most sensitive’) to perturb, or to address the issue of reduced unlearning effect caused by perturbing the parameters of the last few layers. Due to the dependencies among the model’s parameters, we adopt a fine-grained random perturbation unlearning strategy ignoring these dependencies, aiming to explore the differences among the most sensitive unlearning strategy and fine-grained random perturbation strategy. Thus, we propose two strategies for perturbing model parameters, namely Random-k and Top-K.
III-B1 Parameter Selection and Perturbation
@@
For the Top-K strategy, as previously mentioned, it is necessary to evaluate the sensitivity of each model parameter. Certainly, perturbing model parameters one by one and rank their performance change can determine the most sensitive $k$ parameters, but it’s time-consuming. To quickly obtain the sensitivity of parameters, we approximate it using the gradient. Specifically, after the model has converged, yielding the function $F(x,w)$, we introduce a small perturbation to the model parameters denoted as $\epsilon=\{\epsilon_{i,j}\}$. The change in the model’s output can be approximately measured using the following formula:
@@
Where $\bigtriangleup(x,w)$ represents the change in the model’s output performance due to the introduced perturbation $\epsilon$, and $\partial F(x,w)/\partial\theta_{i,j}=g_{i,j}(x)$ is the gradient of learned model to parameter $w$. This evaluation of sensitivity allows us to assess the impact of each parameter on the model’s predictions.
Focusing on aforementioned Equation 4, assuming the perturbation $\epsilon$ added to the model is a same constant for all parameters, the calculation of sensitivity can be simplified as follows:
@@
Fig.2 illustrates the process of our strategies. After completing the parameter screening, we introduce perturbations by drawing upon the insights from [36]. Specifically, we apply a minute amount of noise to the selected parameters as follows:
$w_{i}=w_{i}+\epsilon*w_{i}$
$w_{i}\in\{Top-K(w)\lor{Random-k(w)}\}$ (6)
@@ III-B2 Unlearning Process
For the purpose of eliminating the influence of unlearning data on the model, retraining from scratch, despite consuming substantial computational resources and time, can achieve the optimal unlearning effect. As far as the strategy for machine unlearning is concerned, we desire that the unlearning model $M_{UL}$ approximates the retraining model $M_{RE}$, thus ensuring that the unlearning is moving in the right direction. We aim for both the retraining model and the unlearning model to display similar performance when subjected to $D_{UL}$. However, it clearly goes against the initial intention to perform retraining while conducting unlearning. Despite this, for models trained on i.i.d datasets, the retrained model and $M$ are quite similar, as a small amount of data is unlikely to have a profound impact on the model. Therefore, when calculating the similarity of output distributions, we use $M$ in place of the retrained model in the third step.
Jensen-Shannon (JS) divergence is used to guide the model fine-tuning of the unlearning process [45]. As a symmetric and always finite measure, JS divergence provides a reliable measure of the similarity or distance between two probability distributions. Suppose $x\in D_{UL}$, $P(w,x)$ and $P(\theta,x)$ are posterior distributions of retraining model $M_{RE}$ and unlearning model $M_{UL}$, respectively. $JS$ divergence is calculated as following:
@@
$JS(P(\theta,x)||P(w,x))=-log2+\frac{1}{2}\sum P(\theta,x)\log P(\theta,x)$
$+\frac{1}{2}\sum P(w,x)\log P(w,x)-\log(P(\theta,x)+P(w,x))$ (8)
In the context of machine unlearning, the JS divergence between distributions $P(w,x)$ and $P(\theta,x)$ can quantify the influence of $D_{UL}$ on $M_{UL}$. If the JS divergence is small, the influence of $D_{UL}$ on $M_{UL}$ is deemed to be minor. Therefore, during training, our goal should be to minimize the JS divergence between these distributions to ensure the effectiveness of unlearning. Moreover, given that the value of the JS divergence always falls between 0 and 1, we can intuitively understand the degree of similarity between $M_{UL}$ and $M$.
@@
8:     $w_{j}\leftarrow w_{j}+\epsilon\times w_{j},w_{j}\in\{Top-K(w)\}$
@@
11:     $mask\leftarrow(mask<k)\times\epsilon$
12:     $w\leftarrow(1-mask)\times w+mask\times w_{j},j\in\{Random-k(w)\}$
@@
III-C Unlearning Degree Quantification
By employing a fine-grained model parameter perturbation strategy for the machine unlearning process, the unlearning model and the source model $M$ essentially exhibit the characteristic of model indistinguishability. For datasets with i.i.d attributes, quantifying the unlearning effect by examining the accuracy on the remaining data $D_{RE}$ and the unlearning data $D_{UL}$ of the unlearning model can be challenging. As Fig.1 shows the accuracy of both $D_{RE}$ and $D_{UL}$ may be maintained at a fairly satisfactory level, the difference between these two items in such experiments is about only 15% which is quite small. And Fig.6 displays the same observation, which will be discussed in detail in Section IV-C3.
To address the difficulty in measuring the effect of unlearning and further quantify the degree of forgetting caused by each unlearning strategy, we consider breaking the i.i.d attribute of $D_{UL}$.
@@ III-C Unlearning Degree Quantification
where $\eta$ is a coefficient to balance the $M(\cdot)$’s classification accuracy and $M_{UL}(\cdot)$’s unlearning effect. As Equation.1 explained, our goal is to explicitly demonstrate the degree of unlearning in terms of performance. Once completing the training of SPD-GAN, the performance difference between $M(\cdot)$ and $M_{UL}(\cdot)$ on perturbed data $D_{p}$ is indicative of unlearning degree. That is:
@@
III-D Time Complexity Analysis
@@
We use Equation 5 to approximate the sensitivity of each parameter. Since $|w|=N$, the time complexity for both the forward pass and back-propagation is linear in the number of parameters, the overall time complexity for evaluating sensitivity remains $\mathcal{O}(N)$. Once the model structure is determined, the parameters with Top-K sensitivity of a specific model should be calculated only once. This means that the time required to compute the top-K sensitive parameters will be significantly less than the time taken to perform the unlearning.
IV Experiments
IV-A Baseline Methods
The impact of the first few layers on the model is mainly to extract the features of the data, and the latter few layers use these features to perform tasks. Following the setting of [10], we set $K=\{5,10\}$, use EU-K and CF-K as two of our baseline methods. To track the acceleration of machine unlearning, we also take retraining into consideration.
EU-K: Performing exact unlearning involves targeting the final $K$ layers of the model. In this process, the model parameters are held constant except the last $K$ layers. These last $K$ layers are then reinitialized and trained from the beginning using $D_{RE}$.
CF-K: Catastrophic forgetting the last $K$ layers of the model and then fine-tuning the last $K$ layers using $D_{RE}$. Differing from EU-K, there’s no need to reinitialize the parameters of the last $K$ layers. The parameters of all layers except the last $K$ are held constant.
@@ IV-B Evaluation Metrics
The forgetting rate is a measure of performance decay on the unlearning data, post the model’s unlearning phase. It measures the extent of accuracy decline in the unlearning data after machine unlearning, relative to the accuracy before unlearning.
@@
Here, $Acc_{before}$ refers to the accuracy of model $M$ on the unlearning data before the unlearning process is applied, while $Acc_{after}$ denotes the accuracy of model $M_{UL}$ on the same data set after the unlearning process has been completed.
@@
To assess the degree of model indistinguishability, we employ the $similarity=1-JS~{}divergence$ as a metric to quantify the similarity between the unlearning model and the retraining model, where JS divergence is defined in Eq.8.
IV-C Experimental Results
IV-C1 Training settings
We conduct experiments on CIFAR-10 [46] datasets using ResNet18 [47], VGG [48], GoogLeNet [49], and DenseNet [50] models. The training dataset is randomly divided into $D_{UL}$ and $D_{RE}$ with $D_{UL}$ comprising 5%, 10%, 15% and 20%. For Top-K, we set $K=45$ for ResNet18, and for the Random-k, $k=5\%$. The coefficient $\lambda$ in Equation.9 is set to 0.1. We use Adam as optimizer, for EU-K and CF-K, we set learning rate 1e-4. The generator architecture in SPD-GAN consists of (two Conv2d)-BatchNorm-Conv2d-BatchNorm layers for the encoder, followed by ConvTranspose2d-BatchNorm-ConvTranspose2d-Conv2d layers for the decoder. The LeakyReLU and ReLU activation functions are used for the encoder and decoder, respectively. Activation function Tanh is used to limit the output within [-1, 1]. The coefficient $\eta$ in Equation.12 is 0.03. All the experiments are conducted on a Nvidia A100 GPU.
IV-C2 Determination of $K$ value for Top-K
@@
(a) ResNet18
(b) VGG
(c) DenseNet
(d) GoogLeNet
We aim to achieve the effect of inexact machine unlearning by minimal number of parameter perturbed, to investigate the optimal value of $K$ for Top-K unlearning strategy, we conduct experiments with different values of $K$ within the range of 0 to 50 when the unlearning data ratio varies from 5% to 20%. Following the above settings, we apply Top-K to four models over 100 epochs. Since machine unlearning aims to eliminate the influence of unlearning data while trying to not hurt the performance of remaining data, we calculate the accuracy difference between $Acc_{RE}$ and $Acc_{UL}$. As illustrated in Fig.5, the $K$ value corresponding to the maximum accuracy difference is identified as the optimal. As shown in Tab.I, considering the optimal $K$ values under four unlearning data ratios, we set $K=45$ for ResNet18 and DenseNet, $K=30$ for VGG, and $K=50$ for GoogLeNet.
As shown in Fig.5, although we have chosen the $K$ corresponding to the maximum accuracy difference, in actual experiments, the selection of $K$ values fluctuating within the range of 0 to 50 does not bring significant variability to the Top-K unlearning. For example, in Fig.5(a), for ResNet18, setting K=25 versus 45 does not make a substantial difference. Additionally, measuring the difference in accuracy is a sufficient but not necessary condition for the effectiveness of machine unlearning. Therefore, additional efforts or computational costs to determine a very precise $K$ value are unnecessary; we can roughly set it to a value such as 45.
IV-C3 Effectiveness of Top-K and Random-k
Fig.6 illustrates the unlearning accuracy $Acc_{UL}$ and remaining accuracy $Acc_{RE}$ of four different models trained on the CIFAR-10 using Top-K and Random-k over 50 epochs with unlearning ratio 20%. At the beginning of fine-tuning, in terms of accuracy, the $D_{RE}$ corresponding to Top-K is significantly higher than that of Random-k, indicating that fewer parameter perturbations result in a better memory retention effect. Conversely, the $D_{UL}$ corresponding to Random-k is significantly lower than that of Top-K, suggesting that more parameter perturbations can enhance the extent to which the model’s classification ability on $D_{UL}$ is disrupted.
Fig.1 further shows the accuracy after applying the Random-k on the CIFAR-10 for 50 epochs. The accuracy of $D_{RE}$ is significantly higher than that of $D_{UL}$. It can be observed that the application of the Top-K achieves a 95% $Acc_{RE}$ across the four models, indicating a strong memory retention capability. Meanwhile, the accuracy difference between unlearning model on the $D_{RE}$ and the $D_{UL}$ (both in Fig.6 and Fig.1) shows that both the Top-K and Random-k achieve the unlearning effect more or less.
TABLE I: The number of parameters perturbed in different models under various unlearning strategies.
@@ IV-C3 Effectiveness of Top-K and Random-k
TABLE II: The experimental results of the ResNet18 under different unlearning ratios of the CIFAR-10. Lower $Acc_{UL}(\%)$ and Time (s) and higher other metrics (%) indicate better performance.
$D_{UL}/D$ Strategy $Acc_{UL}$ $Acc_{RE}$ FR MRR Similarity Unlearn Time
5% Top-K 80.01% 97.61% 19.97% 97.63% 88.37% 76.97s
Random-k 76.56% 90.07% 23.44% 90.07% 78.77% 135.43s
EU-5 85.68% 98.44% 14.31% 98.44% 84.88% 534.76s
EU-10 85.88% 99.02% 14.12% 99.02% 69.42% 662.34s
CF-5 86.52% 99.16% 13.48% 99.16% 78.65% 586.6s
CF-10 86.64% 99.54% 13.36% 99.54% 99.98% 607.24s
Retrain 84.84% 100% 15.16% 100% 100% 1190.15s
10% Top-K 80.14% 97.19% 19.86% 97.19% 85.36% 192.86s
Random-k 81.48% 96.52% 18.52% 96.52% 91.03% 249.11s
EU-5 86.7% 98.88% 13.29% 98.88% 82.25% 541.38s
EU-10 84.88% 98.12% 15.12% 98.12% 90.4% 417.87s
CF-5 85.52% 97.89% 14.48% 97.89% 70.27% 515.13s
CF-10 86.26% 98.91% 13.73% 98.91% 91.37% 664.39s
Retrain 84.8% 100% 15.2% 100% 100% 1040.08s
15% Top-K 75.48% 96.8% 24.52% 96.8% 89.88% 203.23s
Random-k 79.85% 98.09% 20.15% 98.09% 83.59% 335.87s
EU-5 84.61% 97.46% 15.38% 97.46% 85.11% 494.60s
EU-10 86.52% 98.75% 13.48% 98.75% 91.38% 627.94s
CF-5 87.64% 99.46% 12.35% 99.46% 85.27% 558.31s
CF-10 86.05% 98.79% 13.94% 98.79% 91.75% 685.18s
Retrain 84.85% 100% 15.15% 100% 100% 950.51s
20% Top-K 75.46% 97.07% 24.54% 97.07% 86.37% 425.01s
Random-k 77.83% 98.99% 22.17% 98.99% 94.51% 638.63s
EU-5 84.48% 97.91% 15.51% 97.9% 71.77% 636.77s
EU-10 83.19% 96.81% 16.81% 96.81% 83.92% 741.51s
CF-5 85.86% 98.14% 14.14% 98.14% 81.76% 778.28s
CF-10 85.69% 99.17% 14.31% 99.17% 91.9% 979.65s
Retrain 85.24% 100% 14.76% 100% 100% 1400.9s
@@
Tab.II provides a more comprehensive overview of the effects of different unlearning strategies at various unlearning ratios. Across different unlearning ratios, Top-K consistently exhibits the lowest $Acc_{UL}$ values (80.14%, 75.48%, and 75.46%), along with corresponding forgetting rates (19.86%, 24.52%, and 24.54%). This suggests that, relative to other strategies, Top-K achieves the most effective reduction of the impact of $D_{UL}$ on the model when the unlearning ratio belongs to 10%, 15%, 20%. However, for the 5% unlearning ratio under Random-k, its corresponding $Acc_{RE}$ and MRR of 90.07% noticeably lag behind the results of other strategies. We believe this disparity is due to the instability of the strategy that employs random parameter perturbation, leading to inconsistent unlearning effects. In the meanwhile, across both $Acc_{RE}$ and MRR, the performance of $M_{UL}$ from nearly all unlearning strategies is comparable to that of retraining models, achieving close to 100% accuracy. As for similarity, there isn’t a significant distinction among different strategies, but this does indicate that model indistinguishability is maintained across diverse strategies.
@@
As discussed earlier, we believe that employing machine unlearning strategies offers a faster way to eliminate the influence of unlearning data on models compared to the retraining from scratch. The metric ’Unlearn Time’ in Tab.II indicates the time required for various strategies to complete unlearning. All strategies achieve an accelerations, but Top-K and Random-k take less time. As illustrated in Tab.I, this acceleration is due to the fact that, for experiments on ResNet18, Top-K and Random-k perturb less parameters than others.
To further illustrate the acceleration effect of unlearning strategies on different models compared to the retraining method, we also conduct machine unlearning tasks on VGG, DenseNet, and GoogLeNet. Fig.7 shows the acceleration ratio of each unlearning strategy compared to retraining under different models and unlearning ratios. The acceleration effect of Top-K and Random-k is better than other strategies. Taking Fig.7(a) as an example, when the unlearning ratio is 5%, Top-K can achieve an acceleration of more than 15x, while Random-k achieves an acceleration close to 9x. The acceleration from EU-K and CF-K methods is not significant. As the unlearning ratio increases, the acceleration of Top-K and Random-k slows down. This is because, as more data is to be forgotten, the degree of modification to the model deepens. In conclusion, we can ascertain that the Top-K achieves optimal unlearning performance with minimal parameter perturbed, maintains considerable memory retention capability, and achieves the fastest acceleration effect.